# Neural ordinary differential equations for model order reduction of stiff systems

### v1

@@ 3 Addressing the stiffness issue
Figure 5: Time reparametrization workflow. Starting from the stiff ODE system (in gray), we compute the discrete solution and obtain the discrete manifold $\mathcal{M}_{h}$ using an implicit method. We normalize the solution, compute the time reparametrization $t_{s}(t)$, and estimate $f_{s}$ (in blue). Finally, we are able to train a nonstiff model that can be solved using an explicit method (in orange).
@@ 3.1 Implementation details
Applying this methodology to the training with the differentiable ODE solver does not require any change. Indeed, the values of the timeseries $\boldsymbol{u}(t_{s})$ (that is, the state space) do not change, they are only differently spaced in time. The main difference concerns the supervised learning approach. As explained before, this step allows us to “initialize” the neural network in a cheap way and requires us to estimate both $\dot{\boldsymbol{u}}(t_{s})=\boldsymbol{f}(t_{s},\boldsymbol{u};\boldsymbol{\mu})$ and $\dot{t}$. The approximation of the derivative is done using a Savitzky-Golay filter with low order (two or three) and small window size ($<10$). Figure 5 shows a scheme of the workflow. To accelerate the training also the values of the time derivative are normalized, namely, we have found a logarithmic normalization to work particularly well.
@@
The neural network $\mathcal{NN}_{t}$ approximating $\dot{t}(\boldsymbol{u}(t_{s};\boldsymbol{\mu}))$ is also trained in a two-step process. First, it is trained in a supervised manner. Similarly to what we do in the other supervised step, we employ as validation loss the accuracy of the predictions on the validation timeseries when employing the ODE solver to integrate both $\dot{t}(\boldsymbol{u}(t_{s};\boldsymbol{\mu}))$ and $\boldsymbol{f}(t_{s},\boldsymbol{u};\boldsymbol{\mu})$. Then, it is fine-tuned using as a target the time $t$. This is achieved by automatic differentiation through Simpson’s integration rule. Namely, the loss is
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{N_{\textnormal{train}}}\sum_{i=1}^{N_{\textnormal{train}}}\left\|t-\int_{\boldsymbol{u}(0)}^{\boldsymbol{u}(t)}\mathcal{NN}_{t}(\boldsymbol{u}(t_{s});\boldsymbol{\theta})\mathrm{d}t_{s}\right\|_{p},$
where the integral is computed cumulatively on the discretization induced by the explicit Runge-Kutta method applied to the trained $\mathcal{NN}$. We tested $p=1,2,4$, the best results were achieved with $p=1$.
@@ 4 Numerical results
Our methodology is tested in five numerical experiments featuring stiff problems widely used in the literature to benchmark implicit ODE integrators [48, 69]. We do not consider dynamical systems derived from semi-discrete PDEs since it would require the usage of autoencoders to extract the latent dynamics. Indeed, we aim to assess the behavior of our method as a stand-alone improvement to neural ODEs. Introducing other variables in the test cases, such as the error of the autoencoders may be detrimental to this task. We are confident that given the wide range of test cases presented, our methodology could be extended to larger state spaces or semi-discrete PDEs without problems. Indeed, being able to outperform classical ODE solvers on small problems without the need to use a simplified latent space representation is a more difficult task.
Computations were carried out on the HPC cluster at MOX with processor ${}^{\textnormal{\tiny\textregistered}}$ ${}^{\textnormal{\tiny\textregistered}}$ Gold 6238R and on a laptop with an ${}^{\textnormal{\tiny\textregistered}}$ ${}^{\textnormal{\tiny\textregistered}}$ 7 PRO 7840U. The reported computational cost refers to the latter CPU. For the sake of reproducibility, we rely on the standard (CPU) implementations found in SciPy [64] of the Runge-Kutta methods. Moreover, to ensure a fair and rigorous comparison with the neural network, we executed both methods on a CPU. While leveraging a GPU would likely enable the neural network to significantly outperform the full-order method, our goal was to maintain a balanced evaluation framework.
@@ 4.1 Test case 1: Van der Pol oscillator
The van der Pol oscillator is a second-order differential equation that exhibits limit cycle behavior, which makes it a valuable model for studying phenomena such as electrical circuits, cardiac rhythms, and neuronal firing patterns. The problem is governed by a parameter $\mu>0$ and we consider its following formulation in the time interval $[0,3.5\mu]$
@@
This test case aims to assess the capabilities of our method on an autonomous periodic system. Indeed, the Van der Pol oscillator has two periodic solutions: a non-trivial periodic solution and an unstable zero solution. The parameter $\mu>0$ determines the significance of the nonlinear component of the equation. When $\mu$ is large the system becomes stiff. We set $\Gamma=[10^{2},10^{4}]$ and consider training dataset composed by a logarithmic discretization $\Gamma_{h}$ of $\Gamma$ with $N_{\mu}=51$ points. The validation dataset is the interval midpoints of $\Gamma_{h}$. The reference solutions are computed using the Radau method, with absolute and relative tolerances set to $10^{-12}$.
@@
The data for the supervised training is smoothed with a Savitzky-Golay filter of window seven and order two. Data, namely the columns of $\boldsymbol{U}$, are randomly subsampled to have about one thousand points for each period of the timeseries. Fine-tuning is done by increasing the size of the unroll length starting from 20 up to 80 (exact values are a hyperparameter to be tuned) and time-stepping $\Delta t=1/40$. The starting learning rate is reduced by a couple of orders of magnitude with respect to the supervised step.
@@
Data is normalized using the following functions: parameter normalization: $\hat{\mu}=\log_{10}\mu$; state normalization: $\hat{\boldsymbol{u}}=(u_{1}/2,u_{2}/\mu)$; dynamics normalization: $\hat{\boldsymbol{f}}=(f_{1}/5,f_{2}/10)/\mu$; time normalization: $\hat{T}=T/\mu$. The model is tested for $N_{\mu}^{\text{test}}=4$ values of $\mu\in[10^{2.01},10^{3.99}]$ spaced logarithmically (these values are chosen so that they are not present in the training nor the validation dataset). Moreover, we are also testing for time beyond what is present in the training and validation datasets, namely, we are integrating until a final time $\hat{T}=5$. Figure 6 shows the results of applying the change of dynamics to this problem. On the left, the plot of the problem with the new dynamics shows smaller gradients. On the right, after the change of variable, we recover with good accuracy the original dynamics of the Van der Pol oscillator. In Table 1 we show a comparison of the computational cost. The number of right-hand side evaluations is one order of magnitude smaller for the explicit solver of the neural ODE w.r.t. the implicit solver of the FOM, this proves that change of variable has significantly reduced the stiffness of the problem. Another interesting feature of our approach is that the ROM error is almost independent from $\mu$, on the other hand, the implicit solver exhibits a correlation among the two, suggesting that for larger values of $\mu$ computational gains may be larger. In this case, the $L^{2}$ error is computed only with respect to the first component of $\boldsymbol{u}$. Indeed, since $u_{2}$ features sharp peaks, small variations in their position make the integral error large and not representative of the accuracy of the solution. For this reason, we have introduced the $d_{\text{peak}}$ metric, which tries to measure the accuracy of the ROM in this regard. We stress that for this test case, the ROM is performing well also for unknown times, maintaining an accurate prediction of the period of the system. From the metrics in Table 1 it is also possible to notice that the main bottleneck in terms of accuracy for the ROM is the time mapping. Indeed, the model is very accurate in the $t_{s}$ parametrization.
@@ 4.2 Test case 2: OREGO problem
$\boldsymbol{\mu}_{\text{test}}\in\{52.5,97.5\}\times\{0.025,0.175\}\times\{10^{-5.975},10^{-4.525}\}.$
Figure 7 shows that the ROM is able to accurately follow the reference solution and track its period. A quantitative comparison with the Radau solver is reported in Table 2. We remark that a tolerance of $10^{-2}$ for the Radau solver did not produce a positive solution for all the problems in the test dataset, thus the comparison is made by employing a tolerance of $10^{-3}$. The number of right-hand side evaluations of the explicit solver is less than half the one of the implicit one, showing that our method has made the system nonstiff. Also, the computational cost is consistently smaller than the one needed by the Radau solver, however, the latter has better accuracy. On the other hand, the distance of the peaks is comparable, proving that the ROM accurately captures the period of the system.
@@
Table 2: Test case 2: OREGO. Comparison of computational cost and accuracy for the Radau solver and the neural ODE based reduced order model (ROM) on the test dataset. Refer to Section 4 for the definition of the metrics.
@@ 4.3 Test case 3: ROBER problem
The system describes the kinetics of an autocatalytic chemical reaction and is considered one of the most popular benchmarks for stiff solvers. When used in this context, the time integration window used is usually large, namely $t\in[0,10^{11})$ and the parameters are $\boldsymbol{\mu}=[0.04,10^{4},3\cdot 10^{7}]$. Indeed, codes might fail if $u_{2}$ accidentally becomes negative, since it then tends to $\infty$- ∞, causing overflow. This feature makes the ROBER problem ideal for testing the ability of our ROM to be accurate across different time scales.
We set $\Gamma=[0.005,0.05]\times[10^{3},10^{5}]\times\{3\cdot 10^{7}\}$ and consider training dataset composed by a logarithmic discretization $\Gamma_{h}$ of $\Gamma$ with 16 uniform points and 31 logarithmically spaced points in the first two directions, respectively. The validation dataset is the interval midpoints of $\Gamma_{h}$. The reference solutions are computed using the Radau method, with absolute and relative tolerances set to rtol$=10^{-10}$, atol$=10^{-14}$.
@@
$\boldsymbol{\mu}_{\text{test}}\in\{0.006,0.49\}\times\{10^{3.025},10^{4.975}\}.$
Figure 8 shows that the ROM is able to accurately follow the reference solution across the different time scales. A quantitative comparison with the Radau solver is reported in Table 3. We remark that a tolerance of rtol$=10^{-4}$, atol$=10^{-7}$ for the Radau solver is among the smallest pair of tolerances that produce a positive solution for all the problems in the test dataset. The number of right-hand side evaluations of the explicit solver is almost one order of magnitude smaller than the implicit one, showing that our method has made the system nonstiff. Moreover, the computational cost is consistently smaller than the one needed by the Radau solver. However, the latter achieves a far smaller error. We highlight that the ROM sometimes fails to reach the correct final time, namely, the map to $t$ sometimes stops at a time much smaller than $10^{11}$. Nevertheless, the ROM still captures the interesting part of the reaction, having issues only when the solution reaches a plateau.
@@ 4.4 Test case 4: E5 problem
The problem describes a chemical pyrolysis where $\boldsymbol{u}$ represents the concentration of the reactants. The largely different rates of reaction that occur in the same system are the cause of stiffness. As a test problem, it is usually integrated in a large time window, namely, we choose the interval $t\in[0,10^{11})$. This means the problem features widely different scales in time and space. The problem is particularly challenging. Indeed, the formulation (12) is affected by the cancellation of digits and it is recommended to use the relation $\dot{u}_{3}=\dot{u}_{2}-\dot{u}_{4}$ to compute $\boldsymbol{f}(t,\boldsymbol{u})$. Moreover, it has been shown that a very small tolerance on the absolute scalar error must be used to obtain a reliable reference solution. Indeed, they are computed using the Radau method, with absolute and relative tolerances set to atol$=10^{-24}$, rtol$=10^{-10}$.
We set $\Gamma=[5\cdot 10^{-10},5\cdot 10^{-9}]\times[10^{7},10^{8}]\times\{1.13\cdot 10^{3}\}\times\{10^{6}\}$ which contains the parameter $\boldsymbol{\mu}=(7.89\cdot 10^{-10},1.1\cdot 10^{7},1.13\cdot 10^{3},10^{6})$, which is the usual choice for the benchmark. The training dataset is built by a discretization $\Gamma_{h}$ of $\Gamma$ with 11 logarithmically spaced points in the first two directions. The validation dataset is the interval midpoints of $\Gamma_{h}$. Data is normalized using the following functions: $\hat{\boldsymbol{\mu}}=\log_{10}\mu$; state normalization $\hat{\boldsymbol{u}}=\log_{10}\boldsymbol{u}/10$; dynamics normalization:
@@
$\boldsymbol{\mu}_{\text{test}}\in\{5\cdot 10^{-9.975},5\cdot 10^{-9.025}\}\times\{10^{7.025},10^{7.975}\}\times\{1.13\cdot 10^{3}\}\times\{10^{6}\}.$

### v2

@@ 3 Addressing the stiffness issue
Figure 5: Time reparametrization workflow. Starting from the stiff ODE system (in gray), we compute the discrete solution and obtain the discrete manifold $\widetilde{\mathcal{M}}$ using an implicit method. We normalize the solution, compute the time reparametrization $t_{s}(t)$, and estimate $f_{s}$ (in blue). Finally, we are able to train a nonstiff model that can be solved using an explicit method (in orange).
@@ 3.1 Implementation details
Applying this methodology to the training with the differentiable ODE solver does not require any change. Indeed, the values of the timeseries $\boldsymbol{u}_{s}(n/N;\boldsymbol{\mu})=\boldsymbol{u}(t^{(n)};\boldsymbol{\mu})$ are known, they only have a different time discretization, that we know pointwise. The main difference concerns the supervised learning approach. As explained before, this step allows us to “initialize” the neural network in a cheap way and requires us to estimate both $\dot{\boldsymbol{u}}_{s}(t_{s};\boldsymbol{\mu})=\boldsymbol{f}_{s}(t_{s},\boldsymbol{u}_{s};\boldsymbol{\mu})$ and $\dot{t}$, in order to train two neural network $\mathcal{NN}$ and $\mathcal{NN}_{t}$, respectively. The approximation of the derivative is done using a Savitzky-Golay filter with low order (two or three) and small window size ($<10$). Figure 5 shows a scheme of the workflow. To accelerate the training also the values of the time derivative are normalized, in particular, we have found a logarithmic normalization to work particularly well.
@@
The neural network $\mathcal{NN}_{t}$ approximating $\dot{t}(\boldsymbol{u}_{s}(t_{s};\boldsymbol{\mu}))$ is also trained in a two-step process. First, it is trained in a supervised manner. Similarly to what we do in the other supervised step, we employ as validation loss the accuracy of the predictions on the validation timeseries when employing the ODE solver to integrate both $\mathcal{NN}_{t}$, which approximates $\dot{t}(\boldsymbol{u}_{s}(t_{s};\boldsymbol{\mu}))$, and $\mathcal{NN}$, which approximates $\boldsymbol{f}_{s}(t_{s},\boldsymbol{u}_{s};\boldsymbol{\mu})$. Then, it is fine-tuned using as a target the time $t$. This is achieved by automatic differentiation through Simpson’s integration rule. Namely, the loss is
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{N_{\textnormal{train}}}\sum_{j=1}^{N_{\textnormal{train}}}\left\|t^{(N)}-\int_{\boldsymbol{u}_{s}(t_{s}^{(0)})}^{\boldsymbol{u}_{s}(t_{s}^{(N)})}\mathcal{NN}_{t}(\boldsymbol{u}_{s}(\tau;\boldsymbol{\mu}^{(j)}),\boldsymbol{\mu}^{(j)};\boldsymbol{\theta})\mathrm{d}\tau\right\|_{p},$
where $t^{(0)},t^{(N)}$ are sample times, and $\boldsymbol{\mu}^{(j)}$ are sample parameters used to build the discrete manifold $\widetilde{\mathcal{M}}$ and the integral is computed cumulatively on the discretization induced by the explicit Runge-Kutta method applied to the trained $\mathcal{NN}$. The time integration interval $(t^{(0)},t^{(N)})$ and the sequence of states $\boldsymbol{u}_{s}(\tau;\boldsymbol{\mu}^{(j)})$ needed for the integration can be picked either from the solution manifold or the values obtained by applying the trained $\mathcal{NN}$ to the training dataset. We tested $p=1,2,4$, the best results were achieved with $p=1$.
@@ 4 Numerical results
Our methodology is tested in five numerical experiments featuring stiff problems widely used in the literature to benchmark implicit ODE integrators [54, 79]. We do not consider dynamical systems derived from semi-discrete PDEs since it would require the use of autoencoders to extract the latent dynamics. Indeed, we aim to assess the behavior of our method as a stand-alone improvement to neural ODEs. Introducing other variables in the test cases, such as the error of the autoencoders may be detrimental to this task. We are confident that given the wide range of test cases presented, our methodology could be extended to larger state spaces or semi-discrete PDEs without problems. Indeed, being able to outperform classical ODE solvers on small problems without the need to use a simplified latent space representation is a more difficult task.
Computations were carried out on a laptop with an ${}^{\textnormal{\tiny\textregistered}}$ ${}^{\textnormal{\tiny\textregistered}}$ 7 PRO 7840U. For the sake of reproducibility, we rely on the standard (CPU) implementations found in SciPy [74] of the Runge-Kutta methods for solving the original stiff system with the original right-hand side. Indeed, the evaluation of the original right-hand side is much cheaper than the evaluation of a neural ODE that approximates it. All the test cases feature systems that cost orders of magnitude more to be solved with explicit solvers than with implicit ones. Moreover, to ensure a fair and rigorous comparison with the neural network, we executed both methods on a CPU. While leveraging a GPU would likely enable the neural network to significantly outperform the full-order method, our goal was to maintain a balanced evaluation framework. We remark that all the presented metrics, tables, and figures refer to the evaluation of $\mathcal{NN}$ and $\mathcal{NN}_{t}$ on a testing dataset, that is with values of the parameter $\boldsymbol{\mu}$ that are unseen by the neural network.
@@
Finally, we indicate with tol the tolerance of the solver. If not otherwise specified, it is intended that we set both the absolute (atol) and relative (rtol) tolerances of the embedded solver to this value.
@@ 4.1 Test case 1: Van der Pol oscillator
The van der Pol oscillator is a second-order differential equation that exhibits limit cycle behavior, which makes it a valuable model for studying phenomena such as electrical circuits, cardiac rhythms, and neuronal firing patterns. The problem is governed by a parameter $\mu>0$ and we consider its following formulation in the time interval $t\in[0,3.5\mu]$
@@
This test case aims to assess the capabilities of our method on an autonomous periodic system. Indeed, the Van der Pol oscillator has two periodic solutions: a non-trivial periodic solution and an unstable zero solution. The parameter $\mu>0$ determines the significance of the nonlinear component of the equation. When $\mu$ is large the system becomes stiff. We set $\Gamma=[10^{2},10^{4}]$ and consider the training dataset composed by a logarithmic discretization $\Gamma$ with $N_{\mu}=51$ points. The validation dataset is the interval midpoints of discretized $\Gamma$. The reference solutions are computed using the Radau method, with absolute and relative tolerances set to $10^{-12}$.
@@
The data for the supervised training is smoothed with a Savitzky-Golay filter of window seven and order two. Data is randomly subsampled to have about one thousand points for each period of the timeseries. Fine-tuning is done by increasing the size of the unroll length starting from 20 up to 80 (exact values are a hyperparameter to be tuned) and time-stepping $\Delta t=1/40$. The starting learning rate is reduced by a couple of orders of magnitude with respect to the supervised step.
@@
Data is normalized using the following functions: parameter normalization: $\hat{\mu}=\log_{10}\mu$; state normalization: $\hat{\boldsymbol{u}}=(u_{1}/2,u_{2}/\mu)$; dynamics normalization: $\hat{\boldsymbol{f}}=(f_{1}/5,f_{2}/10)/\mu$; time normalization: $\hat{T}=T/\mu$. The model is tested for $N_{\mu}^{\text{test}}=4$ values of $\mu\in[10^{2.01},10^{3.99}]$ spaced logarithmically (these values are chosen so that they are not present in the training nor the validation dataset). Moreover, we are also testing for time beyond what is present in the training and validation datasets, namely, we are integrating until a final time $\hat{T}=5$. Figure 6 shows the results of applying the change of dynamics to this problem. On the left, the plot of the problem with the new dynamics shows smaller gradients. On the right, after the change of variable, we recover with good accuracy the original dynamics of the Van der Pol oscillator. In Table 1 we show a comparison of the computational cost. The number of right-hand side evaluations is one order of magnitude smaller for the explicit solver of the neural ODE w.r.t. the implicit solver of the FOM, this proves that the change of variables has significantly reduced the stiffness of the problem. Another interesting feature of our approach is that the ROM error is almost independent from $\mu$, on the other hand, the implicit solver exhibits a correlation among the two, suggesting that for larger values of $\mu$ computational gains may be larger. In this case, the $L^{2}$ error is computed only with respect to the first component of $\boldsymbol{u}$. Indeed, since $u_{2}$ features sharp peaks, small variations in their position make the integral error large and not representative of the accuracy of the solution. For this reason, we have introduced the $d_{\text{peak}}$ metric, which tries to measure the accuracy of the ROM in this regard. We stress that for this test case, the ROM is performing well also for unknown times, maintaining an accurate prediction of the period of the system. From the metrics in Table 1 it is also possible to notice that the main bottleneck in terms of accuracy for the ROM is the time mapping. Indeed, the model is very accurate in the $t_{s}$ parametrization.
@@ 4.2 Test case 2: OREGO problem
$\boldsymbol{\mu}^{\text{test}}\in\{52.5,97.5\}\times\{0.025,0.175\}\times\{10^{-5.975},10^{-4.525}\}.$
Figure 7 shows that the ROM is able to accurately follow the reference solution and track its period. A quantitative comparison with the Radau solver is reported in Table 2. We remark that a tolerance of $10^{-2}$ for the Radau solver did not produce a positive solution for all the problems in the test dataset. When this happens, the error is reported in bold in the table. The number of right-hand side evaluations of the explicit solver is less than half the one of the implicit one, showing that our method has made the system nonstiff. Also, the computational cost is consistently smaller than the one needed by the Radau solver, however, the latter has better accuracy. On the other hand, the distance of the peaks is comparable, proving that the ROM accurately captures the period of the system. In Figure 8 we present a comparison of the stiffness ratio for the original system and the one with simplified dynamics. On average, the time reparametrization reduces the stiffness by a couple of orders of magnitude, showing the effectiveness of our technique.
Figure 8: Comparison of the stiffness ratio $S$ of OREGO system in the original and simplified dynamics.
@@
$10^{-1.5}$ 0.066 3635 196 712 0.0101 0.992
Table 2: Test case 2: OREGO. Comparison of computational cost and accuracy for the Radau solver and the neural ODE based reduced order model (ROM) on the test dataset. Refer to Section 4 for the definition of the metrics. In bold, the error when the solution is not positive.
4.2.1 Parametric initial conditions
  Refer to caption Refer to caption
Figure 9: Test case 2: OREGO, parametric initial conditions. On the left, in colors, the neural ODE predictions in $t_{s}$. On the right, in colors, the prediction is mapped to the original time $t$. In black is the reference solution. The state scale is logarithmic.
To prove that the parameter $\boldsymbol{\mu}$ is not limited to parametrizing the physical properties of the system, in this section we deal with the following variation of Eq. (10)
$\begin{cases}\dot{u}_{1}=77.27(u_{2}-u_{1}u_{2}+u_{1}-8.375\cdot 10^{-6}u_{1}^{2}),\\ \dot{u}_{2}=77.27^{-1}(-u_{2}-u_{1}u_{2}+u_{3}),\\ \dot{u}_{3}=0.161(u_{1}-u_{3}),\\ \boldsymbol{u}(0)=(\mu_{1},\mu_{2},\mu_{3}),\\ \end{cases}$ (11)
where $\boldsymbol{\mu}$ parameterizes the initial condition $\boldsymbol{u}_{0}$. The training samples for the $\boldsymbol{\mu}$ are in
$\Gamma=\{(1,2,3),(5,10,15),(10,20,30)\},$
the normalization of the state space and the right-hand size is the same as before. The parameter $\boldsymbol{\mu}$ is not normalized. The problem is tested on $\boldsymbol{\mu}^{\text{test}}\in\{(2.5,5,7.5),(7.5,15,22.5)\}$. The results are reported in Figure 9. Our model is able to capture the period of the system and the different number of peaks present in the considered time window. The $L^{2}$-error is 0.0408 and 0.0256 in the first and second case, respectively.
@@ 4.3 Test case 3: ROBER problem
The system describes the kinetics of an autocatalytic chemical reaction and is considered one of the most popular benchmarks for stiff solvers. When used in this context, the time integration window used is usually large, namely $t\in[0,10^{11}]$ and the parameters are $\boldsymbol{\mu}=[0.04,10^{4},3\cdot 10^{7}]$. Indeed, codes might fail if $u_{2}$ accidentally becomes negative, since it then tends to $\infty$- ∞, causing overflow. This feature makes the ROBER problem ideal for testing the ability of our ROM to be accurate across different time scales.
We set $\Gamma=[0.005,0.05]\times[10^{3},10^{5}]\times\{3\cdot 10^{7}\}$ and consider training dataset composed by a logarithmic discretization $\Gamma$ with 16 uniform points and 31 logarithmically spaced points in the first two directions, respectively. The validation dataset is the interval midpoints of the discretized $\Gamma$. The reference solutions are computed using the Radau method, with absolute and relative tolerances set to rtol$=10^{-10}$, atol$=10^{-14}$.
@@
$\boldsymbol{\mu}^{\text{test}}\in\{0.006,0.49\}\times\{10^{3.025},10^{4.975}\}.$
Figure 10 shows that the ROM is able to accurately follow the reference solution across the different time scales. A quantitative comparison with the Radau solver is reported in Table 3. We remark that a tolerance of rtol$=10^{-4}$, atol$=10^{-7}$ for the Radau solver is among the smallest pair of tolerances that produce a positive solution for all the problems in the test dataset. The number of right-hand side evaluations of the explicit solver is almost one order of magnitude smaller than the implicit one, showing that our method has made the system non-stiff. Moreover, the computational cost is consistently smaller than the one needed by the Radau solver. However, the latter achieves a far smaller error. We highlight that the ROM sometimes fails to reach the correct final time, namely, the map to $t$ sometimes stops at a time much smaller than $10^{11}$. Nevertheless, the ROM still captures the interesting part of the reaction, having issues only when the solution reaches a plateau. Similarly, we observe that the model is not accurate for $t<10^{-3}$ as it fails to follow the inflection point at $t\sim 10^{-4}$ (cf. the right column of Figure 10). We show that these problems can be mitigated by reducing the tolerance of the RK45 solver to tol$=2\cdot 10^{-5}$, as shown in Figure 11.
@@
Figure 11: Test case 3: ROBER. The prediction of the model with an integration tolerance of tol$=2\cdot 10^{-5}$ mapped to the original time $t$. The smaller tolerance results in more accurate predictions, see Figure 10.
@@ 4.4 Test case 4: E5 problem
The problem describes a chemical pyrolysis where $\boldsymbol{u}$ represents the concentration of the reactants. The largely different rates of reaction that occur in the same system are the cause of stiffness. As a test problem, it is usually integrated in a large time window, namely, we choose the interval $t\in[0,10^{11}]$. This means the problem features widely different scales in time and space. The problem is particularly challenging. Indeed, the formulation (13) is affected by the cancellation of digits and it is recommended to use the relation $\dot{u}_{3}=\dot{u}_{2}-\dot{u}_{4}$ to compute $\boldsymbol{f}(t,\boldsymbol{u})$. Moreover, it has been shown that a very small tolerance on the absolute scalar error must be used to obtain a reliable reference solution. Indeed, they are computed using the Radau method, with absolute and relative tolerances set to atol$=10^{-24}$, rtol$=10^{-10}$.
We set $\Gamma=[5\cdot 10^{-10},5\cdot 10^{-9}]\times[10^{7},10^{8}]\times\{1.13\cdot 10^{3}\}\times\{10^{6}\}$ which contains the parameter $\boldsymbol{\mu}=(7.89\cdot 10^{-10},1.1\cdot 10^{7},1.13\cdot 10^{3},10^{6})$, which is the usual choice for the benchmark. The training dataset is built by a discretization $\Gamma$ with 11 logarithmically spaced points in the first two directions. The validation dataset is the interval midpoints of the discretized $\Gamma$. Data is normalized using the following functions: $\hat{\boldsymbol{\mu}}=\log_{10}\mu$; state normalization $\hat{\boldsymbol{u}}=\log_{10}\boldsymbol{u}/10$; dynamics normalization:
@@
$\boldsymbol{\mu}^{\text{test}}\in\{5\cdot 10^{-9.975},5\cdot 10^{-9.025}\}\times\{10^{7.025},10^{7.975}\}\times\{1.13\cdot 10^{3}\}\times\{10^{6}\}.$